2. (8 points) Evaluate the following integrals. Answers without supporting work will re- ceive no credit. (a) (5 points) | <Evaluate the following integrals. Integral x*ln(x^2+3x+2)dx
Answer #1

2) a) Solm. Now, ali (x²+30+2) := winlfa+i) (0+2)) = nellm (x+1] + ln (+2)) = xm (+1) + xln (0+2) Therefore, Jam(x²+3x+2) daand Jan (2+2) dx = 10(+2) Sadri -Stem (In (427) Sve du] du :- 22 m (2+2) - 1 eta 22 da - 22 /n (1x + 2) - 12 / (x2 - 1) + 4 d